OpenStudy (anonymous):

Given that f(x) = 2x + 5 and g(x) = x − 7, solve for f(g(x)) when x = −3.

OpenStudy (zzr0ck3r):

\[g(x) = x-7 \implies g(-3)=-10\\ f(x) =2x+5 \implies f(-10) = -15\\f(g(-3))=f(-10)=?\]

OpenStudy (anonymous):

-8?

OpenStudy (zzr0ck3r):

what is the last thing I said in line 2?

OpenStudy (anonymous):

wait -15?

OpenStudy (zzr0ck3r):

correct, do you see why?

OpenStudy (anonymous):

omg!!! thank you so much!!!